# Hegel and Modern Topology

### v1

@@ Abstract
In this paper we show how some fundamental concepts of modern topology can be understood philosophically in the light of Hegel’s Logic as well how modern topological concepts can provide concrete illustrations of many of the concepts and deductions that Hegel used. This paper can be seen as a continuation of our paper [13] where we argued that the prototypes of many fundamental notions of modern topology were already found in Aristotle’s Physics.
@@
Let us recall the basic trinary structure of Hegel’s Logic. The Understanding seizes a concept as an immediate whole, but then Dialectical reason remembers the complex process of its genesis and its inward relations and tensions (that it, it becomes a category of objects and morphisms). Speculative reason then replies to Dialectical reason that these relations actually flow in both ways rendering the category again a coherent whole - maybe we can view this as realizing the quotient construction in which we find formal inverses to the relations brought forward by Dialectical reason. Or it is like taking an integral of a certain field density over all possible paths between two points. Thus Speculative reason subsumes into bilateral isomorphisms these directed unilateral relations and again brings peace to the concept, then to be seized again by the Understanding at a higher level.
First look at the Logic of Being
@@
The category of sets $Set$ is an exterior, abstract, discrete concept (Hegel’s concept of number seems to be very set theoretic), but the category of simplicial sets $sSet$ represents a greater cohesion between parts and qualitative structure and determination. The morphisms between objects in a simplicially enriched category instead of being a mere sets become a space.
@@
If $C$ is a category with two objects $A$ and $B$ and if $A$ is initial and $B$ is terminal (there is a canonical $t:A\rightarrow B$) then if there is a morphism $f:B\rightarrow A$ then $A$ and $B$ must be isomorphic. This is a simple prototype of the quotient construction. But let $C$ have three objects, initial 00, terminal $1$ and $A$. Then it follows from these hypotheses that the canonical $0\rightarrow 1$ must factor a unique way as $0\rightarrow A\rightarrow 1$. Now $A$ is allowed to have internal symmetries ($hom(A,A)$ need not be $\{id_{A}\}$). And postulating a $1\rightarrow A$ means that ”$A$ has an element”. Thus $A$ is now a Something.
@@
Besides the quotient construction another fundamental construction is that of completion (but completions and limits in general are often defined using quotient constructions). Take the category of open sets of a topological space. Then we can define a category theoretic diagram $U_{1}\rightarrow U_{2}\rightarrow U_{3}\rightarrow....\rightarrow...$ (equivalent to a net, or local basis) which represents a germ at a given point. But, as we already remarked, the limit in the category-theoretic sense does not in general exist. Thus we need to complete the topology by adding infinitesimal neighbourhoods. So if $G(p)$ represents the germ-object and $U$ some neighbourhood we now have canonical morphisms $G(p)\rightarrow U\rightarrow G(p)$ which represents return-to-self.
These germs or infinitesimals have both the nature of discreteness (they are point-like in the sense that for a 2decent” topology two distinct infinitesimals are represented by disjoint open neighbourhoods around their respective points) and continuity. Open sets are obtained from infinitesimals by ”integration”, a kind of synthesis or gluing together of infinitesimals (cf. the Leibniz notation for the integral). Thus the category of Quantity is born. This is also related to the sheafication construction discussed above.
@@
Conclusion: the quest for Notion
@@
Perhaps a key to understanding Hegel’s concept of Notion and the concept of Individual, Particular and Universal and their relationship - is to see how these concepts are embodied in the movements of Spirit in the Phenomenology of Spirit. That is, we view Notion are describing the structure and dynamics of a community of consciousnesses united in different ways into a higher consciousness wherein each individual also is fully identified with this universal consciousness. Although in the first part of Notion Hegel uses the terminology of the formal logic of his day, what he is getting at is much deeper. There is here the mirroring of the microcosm and the macrocosm, the human mind itself is like a community of minds. This suggests the theory of fractals as well as Alain Connes’ non-commutative geometry - for these all concern interface between the local and global. There are two extremes: the global contained fully in the local (as in holomorphic functions) or the local completely emptied of any global information: this is the case of Penrose tilings which cannot be distinguished by considering only local information. Hegel in discussing Objective Notion mentions a system in which the whole is implicit in each particular determination: this is the case of holomorphic functions: the germ at each point completely determines the whole function (on a connected component containing that point) by analytic continutation and the whole completely determines the germ at each point. In general this is a case of a sheaf in which the restriction morphisms are isomorphisms. Holomorphic functions can be called holistic because they are uniquely determined by their germs, the whole is in the part. Change a holomorphic function locally and you change it globally. The global information on the function is contained implicitly in any local restriction or germ (which unfolds via analytic continuation). Thus we can say that the ”same” germ is spread out across the whole domain of the function. The case of holomorphic function corresponds to the global being contained in the particular (or local). The global is thus totally local. But there is also the opposite situation in which the local is emptied of any global information. The global is totally non-local. This is the case, as we mentioned above, for the space of Penrose tilings. It is impossible to distinguish between two tilings by looking at local patterns. This could also be brought into connection to non-locality in quantum theory.

### v6

@@ Abstract
In this paper we sketch how some fundamental concepts of modern topology (as well as logic and category theory) can be understood philosophically in the light of Hegel’s Science Logic as well how modern topological concepts can provide concrete illustrations of many of the concepts and deductions that Hegel used. Also these modern concepts can in turn be very powerful hermeneutic tools permitting a more rigorous and thorough grasp of Hegelian concepts. This paper can be seen as a continuation of our paper [14] where we argued that the prototypes of many fundamental notions of modern topology were already found in Aristotle’s Physics. More generally it is hoped that this note makes a case for the possibility of a rigorous enriching interaction and mutual support between philosophy on one hand and modern logic and mathematics on the other. This paper is obviously meant only as a preliminary sketch and to offer some motivation for exploring in a more detailed and thorough way the subjects discussed. (January 10th: added section Essence and Metaphysics and section Dialectic of the Combination of Systems).
@@
Also we read in the Lessons in the History of Philosophy:
The two formal moments in this sceptical culture are firstly the power of consciousness to go back from itself, and to take as its object the whole that is present, itself and its operation included. The second moment is to grasp the form in which a proposition, with whose content our consciousness is in any way occupied, exists. An undeveloped consciousness, on the other hand, usually knows nothing of what is present in addition to the content.
@@
Let us recall the basic trinary structure of Hegel’s Logic (see [5] for a deeper discussion). The Understanding seizes a concept as an immediate whole, but then Dialectical reason remembers the complex process of its genesis and its inward relations and tensions (that it, it becomes a category of objects and morphisms). Speculative reason then replies to Dialectical reason that these relations actually flow in both ways rendering the category again a coherent whole - maybe we can view this as realizing the quotient construction in which we find formal inverses to the relations brought forward by Dialectical reason. Or it is like taking an integral of a certain field density over all possible paths between two points. Thus Speculative reason subsumes into bilateral isomorphisms these directed unilateral relations and again brings peace to the concept, then to be seized again by the Understanding at a higher level.
In sections 2-9 we only consider elementary mathematics while in the remaining sections we consider advanced category theory.
First overview of the Logic of Being
@@
Another overview of the Logic of Being
The something and the other. In the Science of Logic Hegel thinks of the idea of two things bearing the same relation to each other and indistinguishable in any other way. Now a good illustration of this situation in found in the two possible orientations of a vector space. Each bears the same relation to the other and there is absolutely no way to uniquely identify one of them in distinction to the other. In the same place Hegel talks about the meaning of proper names and states that they have none !
Hegel’s strategy. All of Hegel is based on the dynamics and structure of consciousness as it reveals itself to itself. But equally important it is all based on positing this structure of consciousness to be essentially objective and not subjective. Thus ’being a subject’ is seen as a stage in the development of the object while mere ’subjectivity’ is seen as a failure and partiality of objective consciousness in not living up to its full objectivity or actuality (or knowledge and realization thereof).
Thus ideality and infinity are the key structure of an object which has developed to a point of being a subject.
Finitude and limit: a closed set. Yet the boundary also defines what the set is not (cf. discussion on Heyting algebra of open sets, etc.). Thus the boundary contains implicitly its own negation.
Limitation : an open set (also analytic continuation). Limit outside itself.
The ought… manifestation of the germ of a space through a given open set representation of the equivalence class. Any particular one is insufficient and can be replaced by another which is also insufficient.
False infinity: taking simply the set (or diagram) of such open set representatives. True infinity taking the limit (equivalence class) or limit (in the categorical sense)..
Degrees of interpenetration and transparency between the individual and the universal (and between self-consciousness and essence) in the Phenomenology of Spirit. From the rudimentary form in the ethical life to: the universal is in the individual, the individual in the universal and the universal is in the relation between individuals and the relation between individuals is in the universal (mutual confession and forgiveness).
Milne’s notes on Class Field Theory has a quote of Chevalley (1940) which seems couched in Hegelian terms: the Abelian extensions of a number field are determined by the number field itself, the number field ’contains within itself the elements of its own transcendence’.
And yet : and yet it is really basic to use ontology and logic inspired by (or abstracted from) biology and psychology - that is, a kind of general systems theory and universal logic based fundamentally on biology and consciousness. The fact that mathematics exhibits philosophical analogy with biology suggests that biological ontology and logic is fundamental. René Thom already did this by suggesting that concepts have analogies with living organisms.
’Pure’ category theory is very poor in results, like Hegel’s poor, empty, abstract universal. But once we consider the particular, like Abelian or Monoidal categories, things become very rich.
Extensive magnitude: the category including in itself objects and arrows. Intensive magnitude, one object signaled out, hence only meaningful in the context of a category. Infinite progression: a diagram in a category. Ratio the limit. For infinitesimals the limit for the category of open sets does not exist but for the functor it does. Ratio is also the quotient construction in which representatives are sublated moments.
@@
The category of sets $Set$ is an exterior, abstract, discrete concept (Hegel’s concept of number seems to be very set theoretic), but the category of simplicial sets $sSet$ represents a greater cohesion between parts and qualitative structure and determination. The morphisms between objects in a simplicially enriched category instead of being a mere set become a space.
@@
If $C$ is a category with two objects $A$ and $B$ and if $A$ is initial and $B$ is terminal (there is a canonical $t:A\rightarrow B$) then if there is a morphism $f:B\rightarrow A$ then $A$ and $B$ must be isomorphic. This is a simple prototype of the quotient construction. But let $C$ have three objects, initial 0, terminal $1$ and $A$. Then it follows from these hypotheses that the canonical $0\rightarrow 1$ must factor a unique way as $0\rightarrow A\rightarrow 1$. Now $A$ is allowed to have internal symmetries ($hom(A,A)$ need not be $\{id_{A}\}$). And postulating a $1\rightarrow A$ means that ”$A$ has an element”. Thus $A$ is now a Something.
@@
Besides the quotient construction another fundamental construction is that of completion (but completions and limits in general are often defined using quotient constructions). Take the category of open sets of a topological space. Then we can define a category theoretic diagram $U_{1}\rightarrow U_{2}\rightarrow U_{3}\rightarrow.$$\rightarrow...$ (equivalent to a net, or local basis) which represents a germ at a given point. But, as we already remarked, the limit in the category-theoretic sense does not in general exist. Thus we need to complete the topology by adding infinitesimal neighbourhoods. So if $G(p)$ represents the germ-object and $U$ some neighbourhood we now have canonical morphisms $G(p)\rightarrow U\rightarrow G(p)$ which represents return-to-self.
These germs or infinitesimals have both the nature of discreteness (they are point-like in the sense that for a ”decent” topology two distinct infinitesimals are represented by disjoint open neighbourhoods around their respective points) and continuity. Open sets are obtained from infinitesimals by ”integration”, a kind of synthesis or gluing together of infinitesimals (cf. the Leibniz notation for the integral). Thus the category of Quantity is born. This is also related to the sheafication construction discussed above.
@@
On Quantity
The indifference of quality which expresses an infinitesimal local quality around a point. The idea of an open set in which the actual size is completely irrelevant. The local homogenous quality. This can be understood in terms of one form of the definition of basis of a topology $\tau$. It is a family of open sets such that each open set $U$ containing a point $x$ contains an element of the basis containing that point. The ordinary topology of $\mathbb{R}^{n}$ has a basis given by open balls of radius $\frac{1}{n}$. The Hegelian ’indifference’ is expressed by the fact that we can remove any finite (or even some infinite sequences) of such balls form the basis without affecting the topology defined by the basis.
The definition of topology is inseparable from the concept of continuity. And yet from the exposition above we see that even the concept of the ordinary topology has an element of discreteness, or rather, can be generated by a discrete sequence.
The concept of continuity, in our opinion, is intimately tied to that of sheaf, in particular with that of gluing or synthesis. In its most rudimentary form it is the condition that any open set must be able to be written as a union $\bigcup_{i\in I}B_{i}$ of open sets in a basis.
Consider a category $C$ of open sets. And let $D:I\rightarrow C$ be a diagram, for instance, balls containing a given point. In this case the point is a ’limit’ of the diagram in some sense, but not in the category $C$. However we can define the notion of $c$-limit.
Given a diagram $D:I\rightarrow C$ for $I$ a directed set, we say that an open set $V$ in $C$ is a $c$-limit of $D$ if i) for all all object $i$ in $I$ we have that $F(i)\cap V\neq 0$ and any open set $V^{\prime}$ satisfying this condition is contained in $V$. For example, for a sequence of balls centered at the origin the complement of the origin is the $c$-limit.
Separation and attraction. Recall the separation axioms in general topology. And the concept of connectivity is also directly involved with separation. The extremes of the coarse and discrete topology. In the coarse topology everything is merged together, we cannot separate things. In the discrete topology not only can we separate everything but each point is its own separation.
Hegel’s ’units’ are like the proximate species of a genus. In Aristotle genus is a continuum, like matter. The species carve up the genus into pieces (which themselves become further genera). But all species form a spectrum, a continuum in sharing the quality of the proximate genus. We can think of a connected groupoid of connected groupoids. Hegel’s unit is like a object in a category. The category of open sets we discussed above. All open sets (at least basic ones) should share the same quality. The category, the topology, is the quality.
Thus we should view quantity as a Kantian process of subjective synthesis - which we already interpreted in terms of the sheaf-gluing process. In the sheafication process we see that a section is essential a unity, a gluing of local section.
Hegel’s treatment of number in the section of Quantum is curious. Indeed given a non-empty finite set $S$ for every $s\in S$ we have $S=S^{\prime}\cup\{s\}$ with $S^{\prime}\subset S$. This is what Hegel seems to call $s$ being a limit of $S$. It matters not which $s$ is used. Hegel’s whole approach suggests an intuitive continuum-based phenomenological view of number, numbers as connected groupoids which expresses the internal homogeneity and indifference between objects.
Different total orderings of a finite set equivalent. For each $S=\{n\in N:n\leq maxS\}$. There are external, imposed. Also external relation of Frege’s equinumerous relation. A number is the set of all sets equinumerous to a given set (the extension of the number ?) … and again this set is irrelevant. An equivalence class. A groupoid.
Hegel is certainly prescient about the relationship between geometry and arithmetic, the discrete and the continuous. Modern mathematics has confirmed this. Numbers do not add themselves, they are added from outside. Arithmetical operations are imposed from without. This is just the category theoretic view of algebraic objects in a category or a natural number object !
Hegel seems to have anticipated induction. Intensive quantity, degree: here we have number as the real numbers, as a set with an order. Now limit is exterior.
Further considerations on Quantity and Measure
The rich connections between Hegel and Kant, Fichte, Schelling, Leibniz and Spinoza have been substantially explored. However there are lesser known and perhaps even more philosophically interesting connections to Sextus Empiricus (as evidenced in Hegel’s Lectures on the History of Philosophy and of course in the Phenomenology itself), Hume and Proclus. The same goes for the various currents of Buddhist philosophy (which also have deep connections to Hume and Pyrrhonism).
In particular there are some very interesting connections to the theory of dialectics and the Proclean view of mathematics in his Commentary of the the first book of the Elements.
A major work relevant for studying Quantity is the PhD thesis of F.M. Nolasco[1]. This thesis suggests the central role of the question of the algebraization of the calculus in Hegel’s critique of the mathematics of his time.
Concerning the section on quantity and the theory of infinitesimals in particular, a central role is played by what might be considered the algebraic expression of infinitesimals (and continuous quantity): attempts to capture infinitesimals via sequences, power series and in particular convergent power series. And the same goes for the related concepts of point, local neighborhood, etc. Perhaps the most interesting embodiment of these concepts is found in algebraic geometry. The idea of completion of a local ring relative to its maximal ideal, which yields a finer or smaller kind of locality (expressed for algebraic varieties by a quotient of a power series ring) then just the local ring. In the standard treatment of completion (relative to a topology which comes from a filtration defined by an ideal) the set of Cauchy sequences is given a purely algebraic description (again through category theoretical ’limits’). There is also the theory of associated graded rings which expresses the tangent cone at a singular point of an algebraic variety. Artinian local local rings are the simplest of all rings, they correspond to points (the Hegelian One), or rather, infinitesimal points which yet have structure.
The theory of meadows, which arose in the study of abstract data types expresses the passage of ring theory (as part of Being) to a part of Essence. A meadow expresses the community and sharing of different rings united by a single point of infinity (analogous to compactification) - this is a very elegant expression of the categories of existence and substance.
In a Peircean terms we can associate Being with monadic predicates, Essence with Relations (or diagrams in a category) and Concept with ternary predicates or indeed with higher order logic (maybe higher category theory), for instance involving the relations between an object and a relation.
@@
The Miracle of Double Negation and Essence
The beginning of the section of Essence in the Science of Logic is acknowledged to be difficult. While the shorted version of the Encyclopedia Logic can be seen more down-to-earth, more connected to concrete embodiments of Essence as pertains to the relativity of being, to classical scientific and metaphysical notions (maybe somewhat in the vein of Whitehead’s Process Philosophy), the treatment of the beginning of Essence in the Science of Logic (specially the section on Schein and the determinations of Reflection) is of a very abstract nature, at once illuminating and frustrating, but suggesting many interesting connections to Hume, Pyrrhonism (skepticism is mentioned in the first remark as being superficial with regards to appearance) and Madhyamaka philosophy. It is almost as if Hegel is struggling to articulate a kind of all-encompassing vision of the impermanence, relativity, lack of svabhava, of the world - all of which leads to the idea of something reflected into itself, shining into itself. The section on Schein can be construed as a metaphysical meditation on change (but also on dialectic itself). It is also Hegel’s solution to Fichte’s problem of the independence of the laws of thought. Hegel deduced them all and expounds how they are all co-implicit thanks to his doctrine of absolute negativity and reflection.
The basic idea is that of pure or absolute negativity which can only maintain itself through self-negation. Negativity posits itself through its own self-negation, that is, negativity creates a semblance, a shining, a reflection of being (at once presupposed, for there has to be something to be negated, and posited, for it is also a result) in the process of its own self-negation. We recommend the reading of the classical books by Hibbens, Harris, Noel, Burbidge and Carlson and D.F. Ferrer’s very detailed discussion of Essence and Reflection in his book A Lógica e Realidade em Hegel (not translated into English).
Thus at the heart of Essence is the negation of negation, double negation. But double negation is one of the richest and deepest of fundamental concepts in mathematical logic, theoretical computer science and category theory. In type theory and computer science it turns out to express the idea of control, of a process having memory and its ability to travel through time and restore a previous state: this is at the heart of the doctrine of Essence. Double negation is tied to the intrusion of imperative programming, and surely the idea of command is linked to that of negation. All command expresse negativity (do this ! means equally: do not do that !) and depends of negation.
We now give a brief description of how double negation arises in computer science connected to the notion of control and jumping through time and memory.
From a type-theoretic point of view there is no one single ’classical logic’. Rather there are many different classical-like extensions of minimal (or intuitionistic) logic, each corresponding to adding different classically but not intuitionistically valid axioms (DNE, LEM, Peirce’s law, etc.). But we do not want mere axioms because they have no computational content and cause problems with term normalization.
Historically some functional programming languages started introducing control operators (typical of imperative programs) which can help handle exceptions or optimize programs. The idea is that in evaluating a term when reaching a subterm involving this operator the process can conditionally jump out (abort the rest of the evaluation of the subterm) of the normal evaluation process and ’pretend’ that the subterm had some other value.
The structure of control operators is related to a style of programming called continuation passing style in which certain functions (which correspond to continuations) are always passed as parameters.
It turns out that if we try to type control operators (such as call/cc in Scheme, or try-catch, or abort, etc.) then these inhabit non-intuitionistically provable types.
And it turns out that writing a term in CPS is equivalent to the Kolmogorov double-negative translation of CL (as a type theory with control) into IPC.
Just as a fragment of IPC is isomorphic to CL, so too CPS style terms can be seen as programs with control (because continuators (which are like the state of the stack) are always passed as arguments).
We recommend F. Agestam’s Interpretations of Classical Logic using $\lambda$-calculus.
@@
A fuller understanding and interpretation of the section on Essence and its relation to modern mathematical physics will be attained through thorough analysis of our use of differential models in science as well as the allied concepts of determinism and computability, specially with regards to the possibility of extending these concepts to other more general mathematical frameworks.
Dialectic of the combination of systems
Given two axiomatic-deductive systems S1 and S2, each is an island unto itself, each exists in the world of its own rules, the symbols of which are meaningful only relative to these rules and to the system as a whole. Thus S1 and S2 are neither mutually consistent nor non-consistent - and we can combine them into a new system in a free way. To speak of consistency we must first identify symbols of both systems, we must somehow postulate a common language. And we need a notion of contradiction. But all this only makes sense relative to a third system S3 which acts as an arbiter. If S1 and S2 deduce contradictory things regarding a symbol A, how can we formalize this in S3 ?
By introducing a symbol for S1 and S2 inside the combined system S1++S2, qualifying symbols representing ’from the perspective of S1’ and ’from the perspective of S2’. The introduction of subscripts or symbols for indexing by S1 and S2 is a projection of the system inside itself, a reflection-into-self.
But then S1 and S2 must be qualified by the arbiter system S3. Why is the perspective of S1 relating to A what it is ? (And the same for S2). S1 and S2 share context S3 and the quality of S1 depends on that of S2 and vice-versa. S1 and S2 have become parts of a whole, for-themselves so far as they are for-another. Communication is the outcome of contradiction.
Essence and Metaphysics
Essence becoming actuality in the Science of Logic: beyond different models for a theory, different frames of reference for physics, representations for an algebraic structure, etc. Do we not here have the phenomenological reflection which neither attempts to have empty thought grasp directly its own structure nor is thought loosing itself in the objectified engagement in its action, but rather is the self-reflected awareness of thought in its thinking, a shift of perspective which knows itself in its process? This goes beyond causality, computation and formal logic to inner and infinite spiritual knowledge. Spiritual knowledge is knowledge that what is in consciousness is taken and proceed from its inmost center, a revelation, veiling-unveiling, which must pass to reveal the process: the revelation is re-velation.
Hegel’s treatment of modality in the section of Actuality is very relevant to contemporary discussions on possible worlds and arguments against the world being a sum of independent contingent facts. There are also connections that could be made to modal logic (i.e. the connection between possibility, actuality and necessity and the axiom $\lozenge P\rightarrow\square\lozenge P$).
The considerations on necessity and substance (and indeed the long digressions on Newtonian forces already suggest this) can be interpreted in terms of the solutions of differential equations (or in general smooth vector fields on a manifold). The local existence and uniqueness of a solution for a given initial conditions expresses determinism which (ideally, though not in mathematical reality) extends to a global determinism.
The first part of Schopenhauer’s World as Will and Representation is divided into four books. It would be interesting to establish a correspondence with the Science of Logic. Clearly books 1 and 2 correspond to Being and Essence while books 3 and 4 correspond to Concept. However books 1 and 2 are written from the perspective ’for us’, from the point of view of transcendental reflection. From this perspective books 1 and 2 belong to Essence and their philosophical knowledge expresses with great rigor and detail Hegel’s theory of appearance, existence, substance, necessity, causality and actuality. Or maybe Hegelian substance is much like the spontaneous production and manifestation of the will in Schopenhauer. It goes beyond phenomena and appearance because the will is immanent in its manifestation, totally in each one and beyond any particular one. The will has to manifest. Transcendental reflection, artistic contemplation and spiritual development represent the will’s progressive self-knowledge, return to self, the Hegelian Spirit in the form of the concept.
Category Theory: a category expresses essence. Each object is a different mode whose determination is inseparable from its relations (morphisms) to the whole (all possible determinations). An object is an expression of the category and yet not the category (for there are other objects). At the same time this circumstance of the object not being the category is itself internalized and expressed as the object being itself the sum-total of its relations (morphisms) with all other objects (i.e. with the totality of the category). This expressed also that the object is the category.
The Concept can be seen as arising when we have the cyclic transcendental consciousness of metatheory. That is to say: we realize that transcendental presuppositions are in cyclic dependency: to do logic we need computation, to do computation we need arithmetic and combinatorics, to these we in turn we need computation and logic. The germ of this in the reciprocal causality at the end of Essence (which we have also associated with higher groupoids). And also, as we have discussed in more detail everywhere, the partial (meta)reflection of one formal system by another (or the formal system by itself) is clearly relevant here: it is the idea of the whole being implicit to some extent in the part. For instance in ZF set theory itself we can prove that there are non-equivalent models of ZF.
The quest for Notion
@@
Perhaps a key to understanding Hegel’s concept of Notion and the concept of Individual, Particular and Universal and their relationship - is to see how these concepts are embodied in the movements of Spirit in the Phenomenology of Spirit. That is, we view Notion as describing the structure and dynamics of a community of consciousnesses united in different ways into a higher consciousness wherein each individual also is fully identified with this universal consciousness. Although in the first part of Notion Hegel uses the terminology of the formal logic of his day, what he is getting at is much deeper. There is here the mirroring of the microcosm and the macrocosm, the human mind itself is like a community of minds. This suggests the theory of fractals as well as Alain Connes’ non-commutative geometry - for these all concern interface between the local and global. There are two extremes: the global contained fully in the local (as in holomorphic functions) or the local completely emptied of any global information: this is the case of Penrose tilings which cannot be distinguished by considering only local information. Hegel in discussing Objective Notion mentions a system in which the whole is implicit in each particular determination: this is the case of holomorphic functions: the germ at each point completely determines the whole function (on a connected component containing that point) by analytic continutation and the whole completely determines the germ at each point. In general this is a case of a sheaf in which the restriction morphisms are isomorphisms. Holomorphic functions can be called holistic because they are uniquely determined by their germs, the whole is in the part. Change a holomorphic function locally and you change it globally. The global information on the function is contained implicitly in any local restriction or germ (which unfolds via analytic continuation). Thus we can say that the ”same” germ is spread out across the whole domain of the function. The case of holomorphic functions corresponds to the global being contained in the particular (or local). The global is thus totally local. But there is also the opposite situation in which the local is emptied of any global information. The global is totally non-local. This is the case, as we mentioned above, for the space of Penrose tilings. It is impossible to distinguish between two tilings by looking at local patterns. This could also be brought into connection to non-locality in quantum theory.